• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Boeglin

    Log entry time 15:56:11 on August 5,2006

    Entry number 177612

    keyword=Shift Summary

    8:00 continue on kinematics 3b run 2154
    8:45 beam position varies (up to 1 mm) due to beam tuning to Hall C

    11:40 start to stwitch to kinematics 11

    12:20 Pointing runs: start with carbon optics target, asked for beam
    13:05 - 13:30 4cm dummy, 15cm dummy, tantalum (took 1 hour)
    14:40 move to rate checks, start with LD and radiator out
    14:55 ready to start rate runs


    kinematics 3b runs
    run# evens current
    2154 2M 4uA
    2155 2M 4uA
    2156 2M 4uA
    2157 2M 4uA
    2158 2M 4uA
    2159 2M 4uA
    2160 2M 4uA
    2161 2M 4uA
    2162 2M 4uA Total 20M

    Carbon optics runs:
    2163 700k 15-40uA checked for optimal current,
    some position fluctuations
    2164 Fast Feed Back on. T3 rate about 1.1kHz.

    Dummy runs: 2165 (4cm),2166 (15cm)
    2167 Tantalum

    Rate checks: 2168 LD 8uA 1.7 kHz no radiator (optimal current up to 10uA)
    2169 LD 3uA 2.6 kHz with radiator
    2170 LD 6uA 2.5 kHz with radiator and PS3=2 (optimal)
    2171 LH 10uA 1.8kHz with radiator and PS1=1
    2172 LH 20uA 0.4kHz no radiator and PS1=1

    For production:

    LD+radiator: PS3=2, Current about 8uA check dead time < 25%
    LD alone: PS3=1 , Current 9uA

    LH+radiator: PS3=1, current 10uA
    LH alone: PS3=1, current 20uA